### v1

@@ 6.8.1 Comparison with MASKCRYPT (10506637, )
In contrast, our FAS method provides a more efficient balance between security, performance, and scalability. By combining selective homomorphic encryption with differential noise and bitwise scrambling, our approach encrypts a fixed percentage of weights (e.g., 10%) without the need for gradient sensitivity analysis or pre-training. This allows us to achieve high computational efficiency, with a 90% reduction in processing time compared to full homomorphic encryption, while also providing robust resistance to model inversion attacks. The layered privacy provided by noise and scrambling further enhances security without incurring the extra communication and computational costs associated with MASKCRYPT’s per-round gradient analysis process. Importantly, because our method doesn’t require per-round gradient recalculations or client-specific setups, it scales effectively across larger federated networks without additional setup costs for each client.
6.8.2 Comparison with FedML-HE (jin2024fedmlheefficienthomomorphicencryptionbasedprivacypreserving, )
@@
FedML-HE employs a targeted encryption strategy in federated learning by pre-training the model to identify critical weights based on gradient sensitivity, which are then selectively encrypted.FedML-HE model was recreated by us to evaluate its performance. While the results may not fully match the original models values, they offer a reliable insight into FedML-HE’s behavior. While this approach offers strong privacy protection for the most sensitive weights, it introduces notable time and resource costs due to the pre-training phase. Additionally, in FedML-HE, each client independently generates an encryption mask on its local dataset, which is then aggregated at the central server. While this avoids directly sharing a mask across clients, it can limit scalability, as each client’s mask may vary based on smaller, locally available datasets, which may lead to inconsistencies in aggregated model accuracy.
Our FAS method avoids these limitations by providing privacy without the need for pre-training or mask aggregation. Instead of selecting weights through gradient sensitivity, our method encrypts a fixed percentage of weights and augments this with differential noise and bitwise scrambling. This eliminates the computational burden of FedML-HE’s pre-training phase and simplifies scalability, as our approach does not depend on client-specific training data characteristics. At 10% encryption, our approach achieves comparable security results (e.g., MISSIM score of 58%) to FedML-HE’s pre-trained encryption (52%) but with significantly reduced complexity, better scalability, and faster performance.
In summary, our FAS method provides a more efficient and privacy-preserving alternative to FedML-HE, particularly suited for federated learning environments where computational resources and data confidentiality are key considerations. By reducing the computational burden and ensuring consistency across clients without needing mask aggregation, FAS is easily scalable to large federated networks.
Figure 15: Comparison of MSSIM Scores: Our work vs. FedML-HE vs. MASKCRYPT
Figure 16: Comparison of VIFP Scores: Our work vs. FedML-HEvs. MASKCRYPT
@@
Figures 15 and 16 provide a visual comparison of our FAS method, FedML-HE, and MASKCRYPT. In Figure 15, we observe that at 10% encryption, FedML-HE demonstrates slightly higher stability with a 52% MSSIM score compared to our 58%. MASKCRYPT, on the other hand, achieves a 55% MSSIM score, placing it between the two. However, both FAS and MASKCRYPT converge with FedML-HE around 20% encryption, resulting in nearly identical security performance across all methods. Figure 16 similarly shows that while FedML-HE achieves marginally better stability at lower encryption levels, and MASKCRYPT performs slightly better than FAS in these cases, our FAS technique provides comparable security with significantly reduced computational costs. MASKCRYPT incurs a higher computational overhead due to its sensitivity mask recalculation in every round, making it less efficient in scenarios requiring real-time performance. This efficiency gap makes FAS an ideal choice for scalable and privacy-sensitive applications, particularly in real-time federated learning scenarios.
In conclusion, each method has unique advantages based on application needs. For cases where lower encryption levels suffice and computational time is less critical, FedML-HE or MASKCRYPT may offer a slight edge in stability. However, for environments that demand higher encryption percentages along with performance and scalability, our FAS approach provides equivalent security with reduced time costs. This makes FAS a more efficient and scalable choice for federated learning, particularly in applications that prioritize real-time performance.
Table 7: Comparison of Encryption Techniques for Different Models
@@
Table 7 shows the non-encrypted, partly encrypted, and overhead times for different models using three encryption methods: Our FAS, FedML-HE, and MASKCRYPT. The data in the table is a logical reconstruction of the models based on the methodologies described in the respective papers, and may not exactly represent the original implementations. Our FAS consistently has the lowest overhead across models, making it the most efficient. FedML-HE shows the highest overhead due to pre-training, while MASKCRYPT offers a balance but initially has lower security, catching up after several rounds. Overall, our FAS is the best option for efficiency and scalability in federated learning.
6.9 Comparative Analysis of Encryption Techniques Across Datasets
This section provides a detailed comparison of the three encryption techniques—FAS, FEDML-HE, and MASKCRYPT—across five datasets: CIFAR-10, Diabetic Retinopathy, COVID, Lung, and Kidney datasets. The analysis emphasizes the efficiency of FAS in reducing both overhead and encryption time compared to the other methods.
6.9.1 CIFAR-10 Dataset
Figure 17 demonstrates the performance on the CIFAR-10 dataset. FAS consistently achieves the lowest total training times and overhead across all models. Unlike FEDML-HE, which incurs significant pretraining costs for sensitivity mask creation, and MASKCRYPT, which generates masks at each round, FAS employs a random masking strategy with minimal computational burden. This positions FAS as the fastest and most efficient method for CIFAR-10.
6.9.2 Diabetic Retinopathy Dataset
As depicted in Figure 18, FAS significantly reduces overhead times for the diabetic retinopathy dataset compared to FEDML-HE and MASKCRYPT. The pretraining in FEDML-HE and the per-round mask generation in MASKCRYPT add substantial time costs, making FAS the preferred choice for this dataset.
6.9.3 COVID Dataset
The results for the COVID dataset are presented in Figure 19. FAS demonstrates a clear advantage, particularly in lightweight models like MobileNetV2, where its simplicity translates to much lower overhead times. The benefits of avoiding sensitivity mask-related computations are most apparent in this dataset.
6.9.4 Lung Dataset
Figure 20 showcases the performance on the lung dataset. FAS continues to outperform the other techniques by a significant margin, particularly in larger models such as ResNet-50. This demonstrates its scalability and adaptability to more complex models and datasets.
6.9.5 Kidney Dataset
Finally, Figure 21 highlights the results for the kidney dataset. Across all models, FAS achieves the best performance, confirming its robustness and suitability for diverse datasets. The elimination of pretraining and per-round computations makes FAS a practical and efficient solution.
@@ 6.9.5 Kidney Dataset
Figures 17, 19, 18, 21 and 20 shows that across all datasets, FAS consistently outperforms FEDML-HE and MASKCRYPT in terms of total training time and overhead. The absence of sensitivity mask pretraining or extra per-round computations is the key factor contributing to its superior performance. These results establish FAS as the optimal choice for applications requiring fast and efficient encryption without compromising security. We can see that in smaller datasets which training time is not significant, difference between encryption technique becomes more obvious.
6.10 Handling Data Skew in Encryption-Based Techniques
@@
6.10.1 Challenges with Mask-Based Techniques
@@ 6.10.2 Robustness of the FAS Technique
6.10.3 Experimental Results and Comparisons
Figure 22: MSSIM and VIFP score comparison for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 23: MSSIM and VIFP score comparison for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figure 24: MSSIM and VIFP score comparison on lung dataset for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 25: MSSIM and VIFP score comparison on lung dataset for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figure 26: MSSIM and VIFP score comparison on kidney dataset for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 27: MSSIM and VIFP score comparison on kidney dataset for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figure 28: MSSIM and VIFP score comparison on covid dataset for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 29: MSSIM and VIFP score comparison on covid dataset for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figure 30: MSSIM and VIFP score comparison on diabetic retinopathy dataset for 10% encryption under skewed and normal data conditions. FAS maintains consistent performance, whereas MASKCRYPT and Fedml-HE show degraded results under skewed data.
Figure 31: MSSIM and VIFP score comparison on diabetic retinopathy dataset for 10% encryption under skewed and normal data conditions. Similar trends are observed, with FAS showing robustness to data skew.
Figures 22, 23, 24, 25, 26, 27, 28, 29, 30, 31 shows that the proposed FAS technique, with its reliance on encryption, scrambling, and noise, demonstrates superior robustness to data skew across all evaluated datasets—Kidney, Lung, COVID, and Diabetic Retinopathy—compared to existing gradient-based encryption methods. This property ensures that FAS retains its performance and security benefits even in diverse and challenging real-world scenarios. The results from experiments using the EfficientNetB0 model across these datasets highlight that the accuracy of FAS is minimally impacted by data skew, unlike other methods, which exhibit noticeable performance degradation under similar conditions. This consistent robustness across datasets underscores the adaptability and reliability of the FAS technique for privacy-preserving machine learning.

### v2

@@ VI-H1 Comparison with MASKCRYPT
In contrast, FAS stabilizes earlier by combining selective homomorphic encryption, differential noise, and bitwise scrambling, encrypting a fixed percentage of weights without gradient sensitivity analysis or pre-training. This approach reduces processing time by 90% compared to full homomorphic encryption while maintaining strong resistance to model inversion attacks. Unlike MASKCRYPT, which recalculates sensitivity masks each round, FAS eliminates extra computational costs and scales effectively across federated networks.
VI-H2 Comparison with FedML-HE
@@
FedML-HE [18] identifies and encrypts critical weights based on gradient sensitivity through a pre-training phase. Our recreated FedML-HE model provides insight into its behavior, though results may differ from the original implementation. While FedML-HE offers strong privacy protection, its pre-training phase introduces significant time and resource costs. Each client generates an encryption mask independently, which, while avoiding direct mask sharing, can lead to inconsistencies in aggregated model accuracy.
Our FAS method eliminates the need for pre-training and mask aggregation by encrypting a fixed percentage of weights with differential noise and bitwise scrambling. At 10% encryption, FAS achieves a MSSIM score of 58%, comparable to FedML-HE’s 52%, but with significantly lower computational costs. The lack of pre-training and per-client mask generation allows FAS to maintain efficiency and scalability in federated learning environments.
Figure 12: Comparison of MSSIM and VIFP Scores: FAS vs. FedML-HE vs. MASKCRYPT
@@
Figure 12 compares FAS, FedML-HE, and MASKCRYPT across MSSIM and VIFP scores. At 10% encryption, FedML-HE shows slightly higher stability (MSSIM 52%), while FAS achieves 58%, and MASKCRYPT falls in between at 55%. As encryption increases to 20%, all methods converge with nearly identical security performance. Despite slight advantages for FedML-HE in lower encryption levels, FAS provides comparable security at significantly lower computational costs. MASKCRYPT’s need for sensitivity mask recalculation each round makes it less efficient for real-time scenarios. These findings position FAS as an optimal choice for scalable, privacy-sensitive federated learning applications requiring low-latency performance.
TABLE III: Comparison of Encryption Techniques for Different Models
@@
Table III compares non-encrypted, partly encrypted, and overhead times for different models using FAS, FedML-HE, and MASKCRYPT on the Kidney dataset. The values are reconstructed based on methodologies described in the respective papers. FAS consistently has the lowest overhead across models, making it the most efficient. FedML-HE has the highest overhead due to pre-training, while MASKCRYPT offers a balance but initially has lower security, improving after multiple rounds. Overall, FAS is the best option for efficiency and scalability in federated learning.
VI-I Comparative Analysis of Encryption Techniques Across Datasets
This section compares the performance of FAS, FEDML-HE, and MASKCRYPT across five datasets: CIFAR-10, Diabetic Retinopathy, COVID, Lung, and Kidney. The analysis highlights FAS’s efficiency in minimizing overhead and encryption time compared to the other methods.
Across all datasets, FAS consistently achieves the lowest training times and computational overhead as we can see from Figures 13,  14,  17,  16,  15. Unlike FEDML-HE, which incurs significant pretraining costs for sensitivity mask creation, and MASKCRYPT, which generates masks at each round, FAS employs a random masking strategy with minimal computation. This advantage is particularly evident in lightweight models like MobileNetV2 on the COVID dataset and larger models like ResNet-50 on the Lung dataset. The efficiency of FAS is further demonstrated in the Diabetic Retinopathy and Kidney datasets, where it eliminates pretraining and per-round computations, making it the most practical and scalable approach. Overall, FAS outperforms the other techniques across all datasets, confirming its robustness and adaptability.
@@ VI-I Comparative Analysis of Encryption Techniques Across Datasets
Figures 13, 15, 14, 17 and 16 shows that across all datasets, FAS consistently outperforms FEDML-HE and MASKCRYPT in terms of total training time and overhead. The absence of sensitivity mask pretraining or extra per-round computations is the key factor contributing to its superior performance. These results establish FAS as the optimal choice for applications requiring fast and efficient encryption without compromising security. We can see that in smaller datasets which training time is not significant, difference between encryption technique becomes more obvious.
TABLE IV: Top 3 Cases Where FAS Outperforms Other Models
Dataset Model Compared To Enc. Improvement (%) Overhead Improvement (%)
COVID MobileNetV2 FEDML-HE 42.50 73.91
COVID EffNetB0 FEDML-HE 46.15 69.77
COVID EffNetB0 MASKCRYPT 36.36 60.61
Table IV presents the top three cases where the FAS encryption model demonstrates a higher efficiency compared to FEDML-HE and MASKCRYPT in terms of encryption time and overhead reduction.
COVID dataset with MobileNetV2: FAS achieves a 42.50% faster encryption time and a 73.91% lower overhead than FEDML-HE. COVID dataset with EffNetB0: FAS provides a 46.15% encryption improvement and a 69.77% overhead reduction compared to FEDML-HE. COVID dataset with EffNetB0 (vs. MASKCRYPT): FAS reduces encryption time by 36.36% and overhead by 60.61%. These results suggest that FAS is particularly effective in reducing computational overhead while maintaining faster encryption speeds, especially in the COVID dataset with lightweight models.
VI-J Handling Data Skew in Encryption-Based Techniques
@@
VI-J1 Challenges with Mask-Based Techniques
@@ VI-J2 Robustness of the FAS Technique
VI-J3 Experimental Results and Comparisons
Dataset Method MSSIM (Skew/Normal) VIFP (Skew/Normal)
Kidney Fedml-HE 63 / 57 18 / 13
FAS 62 / 61 16.5 / 15
MASKCRYPT 63 / 60 18 / 15
Control 70 / 70 20 / 20
Lung Fedml-HE 65 / 59 19 / 14
FAS 62 / 61 16.5 / 15
MASKCRYPT 66 / 62 19 / 16
Control 70 / 70 20 / 20
COVID Fedml-HE 66 / 60 20 / 15
FAS 62 / 61 16.5 / 15
MASKCRYPT 67 / 63 20 / 17
Control 70 / 70 20 / 20
Diabetic Retinopathy Fedml-HE 64 / 58 18 / 14
FAS 62 / 61 16.5 / 15
MASKCRYPT 65 / 61 19 / 16
Control 70 / 70 20 / 20
CIFAR-10 Fedml-HE 62 / 58 17 / 13
FAS 61 / 59 16.5 / 15
MASKCRYPT 64 / 59 18 / 14
Control 70 / 70 20 / 20
TABLE V: MSSIM and VIFP Scores for Skew and Normal Data across Datasets
Table V demonstrates that the proposed FAS technique, leveraging encryption, scrambling, and noise, exhibits superior robustness to data skew across all evaluated datasets—Kidney, Lung, COVID, and Diabetic Retinopathy—when compared to existing gradient-based encryption methods. This characteristic ensures that FAS maintains its performance and security advantages even in diverse, real-world scenarios with varying data distributions. Experimental results using the EfficientNetB0 model across these datasets show that the accuracy of FAS is minimally affected by data skew, while other methods experience significant performance degradation under similar conditions. These findings highlight the adaptability and reliability of the FAS technique in privacy-preserving machine learning applications.